the ratio of two numberds is 3:4. If the difference is 12 find its smaller number

Answer by fcabanski(1391)

Call the numbers x and y


x/y = 3/4 and y-x = 12 so y=x+12


Substitute that y value (x+12) into the proportion and then solve for x.


x/(x+12) = 3/4 ---> 4x = 3(x+12) = 3x + 36


x=36


y-36 = 12 ---> y = 48


Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)